OpenStudy (anonymous):

Help me prove LHS is equal to RHS? (sin x/1-cos x) + (sin x/1+cos x) = 2 csc x

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@plohrr

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (mertsj):

Write the two fractions on the left as a single fraction using the common denominator (1-cosx)(1+cosx)

OpenStudy (anonymous):

Distribute.

OpenStudy (anonymous):

So we get sin/1-cos^2?

OpenStudy (mertsj):

Here is what you get: \[\frac{\sin x(1+\cos x)+\sin x(1-\cos x)}{(1-\cos ^2x)}\]

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

Th original was: \[\frac{ \sin x }{ 1-\cos x } + \frac{ \sin x }{ 1 + \cos x} = 2 \csc x\]

OpenStudy (mertsj):

Write these two fractions with a common denominator: \[\frac{3}{7}+\frac{3}{11}=\frac{3(11)+3(7)}{7(11)}\]

OpenStudy (mertsj):

Now apply that principal to your problem

OpenStudy (anonymous):

Oh, okay. I see now. Hang on just a minute. :P

OpenStudy (anonymous):

Does the cos x sin x and -cos x sin x cancel out?

OpenStudy (mertsj):

You now have a 2 term numerator. The two terms have sin x as a common factor. Factor it out.

OpenStudy (anonymous):

I don't get it..

OpenStudy (mertsj):